OpenStudy (anonymous):

Find the missing number in each group. 1) Average is 30. Find the missing number: 14, 28, .....

OpenStudy (anonymous):

14+28=7X6=42 So, (42+x)/2=30 42+x=60 x=18

OpenStudy (stacey):

Since the average is of three numbers, that should be (42+x)/3 =30